OpenStudy (anonymous):

Matt purchased a washer and dryer for $2,885 using an 18 month deferred payment plan with an interest rate of 26.12%. What is the account balance after there deferment period if payments of $85 were made each month?

OpenStudy (tkhunny):

You will need a description of EXACTLY what it is that is deferred. 2885*0.2612/12 = 62.80 Okay, so $85 / month will cover the monthly interest. Where does that leave us?

OpenStudy (anonymous):

Do you divide 85 by 62.80, I got $1,355.00 @tkhunny

OpenStudy (anonymous):

& thank you for replying @tkhunny

OpenStudy (tkhunny):

Why would you do that division? Where did you get $1,355.00?

OpenStudy (anonymous):

Idk I'm quite confused, I just thought that you would divide the two

OpenStudy (tkhunny):

Why? Answer that and you will diminish your confusion.

OpenStudy (tkhunny):

How does the deferment plan work? Is the NO interest during the deferral period? 2885 - 18*85 = 2885 - 1530 = 1355 Is interest accrued during the deferral period, you just not required to pay it? 2885 + 1047.07 - 1530 = 2402.07 Does it SEEM like there is no interest during the deferral period but it magically appears if you don't pay it off in the 18 months? 2885 + 1047.07 - 1530 = 2402.07 What EXACTLY are the terms?